OpenStudy (anonymous):

Sea water conains roughly 28.0g of NaCl per liter. What is the molarity of sodium chloride in sea water ?

OpenStudy (anonymous):

step 1 convert the gram to moles as the molarity is \[\frac{ mol }{ L }\] so the molar mass of NaCl is 58.44 so take \[28.0g NaCl * \frac{ 1 mol }{ 58.44g NaCl }*\frac{ 1 }{ 1 L}\]